# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

### v2

@@ 4.2 Experimental Settings
4.3 Results of Evasion Attacks
Tables 2-4 show the results of evasion attacks in the pre-processing, training, and post-training phases, respectively. We also explore advanced attacks by combining different attacks (details in Appendix F). Our key findings are as follows.
A key finding is that dataset auditing methods relying on the embedding of strong external features ($e.g.$, DVBW and UBW-P) in our experiments exhibit relatively higher robustness. It is particularly exhibited against challenging training-phase attacks like adversarial training and DP-SGD. The two methods are poison-label methods that necessitate changing the labels of part of the samples in the dataset. However, even these methods are not immune to evasion attacks, showing vulnerabilities to specific pre-processing data transformations ($e.g.$, WaveletFilter, AutoEncoder) and post-training inference manipulations ($e.g.$, Reprogramming). Besides, other EF-based methods tested using mild external features ($i.e.$, UBW-C, ZeroMark, DW) displayed considerably less robustness across various attack stages.
Moreover, IF-based methods generally exhibit moderate robustness, with DUA demonstrating relatively higher robustness. DUA enhances auditing effectiveness by injecting additional noise into the dataset. This mechanism improves robustness but also makes the audit results more susceptible to forgery. We will further discuss the details in Section 4.4.
Table 5: Results of forgery attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack FGSM PGD UAP TIFGSM VNIFGSM
IF-based MIA Score ($\downarrow$) 0.54 0.66 0.72 0.54 0.71 0.71
Rapid Score ($\downarrow$) 0.22 0.79 0.89 0.30 0.86 0.88
DI p-value ($\uparrow$) 0.01 $10^{-13}$ $10^{-5}$ $10^{-23}$ 0.97 0.01
Diff ($\downarrow$) 0.79 1.35 0.98 1.61 -0.81 0.09
DUA Cost (logits) (%, $\uparrow$) 100.00 0.13 0.13 0.18 0.13 0.24
Cost (label) (%, $\uparrow$) 100.00 0.13 0.13 0.24 0.13 0.24
EF-based DVBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 10.01 29.92 98.37 65.60 69.77 96.77
UBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 11.73 86.55 100.00 71.08 95.69 100.00
ZeroMark p-value ($\uparrow$) 1.00 1.00 1.00 0.99 0.01 1.00
WSR (%, $\downarrow$) 10.01 29.93 98.44 68.28 71.68 96.88
DW WSR (%, $\downarrow$) 11.73 67.00 89.40 66.37 73.72 89.36
4.4 Results of Forgery Attacks
For forgery attacks, we assume that the adversary wants to make their own dataset ($i.e.$, the attack auxiliary dataset) to be falsely identified as the ‘trained dataset’ of other independent models. We use the model trained on the unmarked release dataset as the independent model. The results of white-box forgery attacks are in Table 5, and the experiments in the black-box scenario can be found in Appendix I.
Figure 3: Relative time overhead of different evasion attacks. The overhead is compared with the baseline time without any specific evasion attack (represented as ‘No Attack’ with a relative overhead normalized to 1.0).
Takeaway 4: Existing dataset auditing methods are overall vulnerable to forgery attacks.
Takeaway 5: Dataset auditing methods that involve more complex decision-making and auditing processes are more resistant to forgery.
4.5 Attack Overhead
4.6 Summary of Evaluated Methods
Our analysis reveals that no single method excels across all five dimensions. The most critical finding is the pervasive lack of adversarial robustness and distinctiveness, as also analyzed in the previous section. On the other hand, fidelity is typically high, indicating minimal impact on model utility in benign settings. A clear divide exists in stealthiness. IF-based methods inherently excel in stealthiness, while EF-based methods depend on the embedded feature. Efficiency spans a wide spectrum, from computationally cheap approaches to highly demanding ones requiring intensive processing.
5 Discussion
5.1 Related Works
Since dataset auditing in DL is an emerging topic, there are only a limited number of surveys or benchmarks that have explicitly explored this domain. Asswad et al.[2] highlighted the importance of protecting data ownership and reviewed representative related techniques. Chandrasekaran et al.[10] provided a comprehensive study on machine learning governance, including data ownership in machine learning. Recently, Du et al. [20] presented the first systematic SoK regarding dataset auditing, primarily focusing on surveying existing solutions. From the broader perspective of AI copyright protection, the majority of existing work focuses on watermarking techniques for models [69], generated content [70, 88], and traditional multimedia [55]. Lukas et al. [58] presented a toolbox to evaluate the robustness of DNN watermarking, which is fundamentally different from dataset auditing.
Compared with existing work, our paper makes the following four distinct contributions to the field. (1) This paper focuses on a specific scenario ($i.e.$, auditing the usage of a dataset) where existing techniques to watermark other objects are not feasible. The dataset is specifically used for training a DL model, and the model is the audit target. (2) This paper aims to provide a novel adversarial angle for evaluating existing dataset auditing methods. (3) Existing works fail to provide a comprehensive benchmark and evaluation regarding the SOTA dataset auditing methods. (4) Existing papers rarely discuss forgery attacks against dataset auditing.
5.2 Dataset Auditing for Generative Models
While dataset auditing has been actively explored for classification models, auditing generative models has only recently started to receive attention [60, 67, 39, 48] and remains largely underdeveloped. Many recent results suggest that even achieving stable results useful for auditing in real-world non-adversarial settings is extremely challenging [22, 23]. This difficulty arises from several inherent characteristics of generative models: (1) They are often trained on massive corpora where any individual dataset constitutes only a minuscule fraction, making both internal and external auditing signals significantly diluted. (2) The outputs of generative models are non-deterministic and vary across sampling seeds, further complicating reliable attribution. (3) Many real-world misuse cases manifest at the output level ($e.g.$, a generated image resembling a copyrighted artwork), where neither internal model access nor query APIs are available ($i.e.$, the box-free setting). In such cases, traditional auditing methods, which rely on model behaviors or access, become ineffective or even inapplicable. While very recent works have achieved partial success in specific settings ($e.g.$, in the fine-tuning scenarios [67, 35, 48]), how to design stable and general-purpose auditing algorithms for generative models is still an open challenge in the community.
Given these challenges, existing auditing methods for generative models are not maturely applicable and may yield unstable or misleading results when evaluated under adversarial settings. Therefore, we caution that rushing into adversarial robustness evaluations in such an ill-defined setting without reliable baselines may misrepresent the true difficulty of the problem or misguide the community toward premature conclusions. Nevertheless, arguably, our framework and attack methodology, structured by three processing stages ($i.e.$, pre-processing, training, post-training) and three attack strategies ($i.e.$, decoupling, removal, detection), are data-agnostic and could guide evaluations across different modalities and tasks. We leave this as an important direction for future work.
5.3 Potential Future Direction
Future Direction for Dataset Auditing. The demonstrated vulnerability of existing dataset auditing methods underscores an urgent need for fundamentally more robust and distinctive techniques. Future work may focus on developing novel auditing mechanisms resilient to these attacks. (1) For evasion, this could involve exploring IFs that are harder to decouple from the model utility. For EF-based methods, future works may seek EFs that are deeply integrated into the model’s core functionality rather than easily removable or detectable trigger patterns. A certified robust mechanism is also urgently needed. (2) For forgery, our empirical study suggests that complicating the auditing process may lead to better resistance. Consequently, exploring more complex methods, such as multi-bit watermarks [75] or cryptography [27], may be a promising future direction to improve dataset auditing.
Limitations and Future Directions for Evaluating Dataset Auditing. While DATABench provides a comprehensive benchmark, there are still some limitations. First, this paper primarily focuses on dataset auditing regarding image datasets since they are currently the most widely studied and feasible methods. There have been some preliminary works studying auditing datasets of other modalities, such as text [42, 84], point clouds [83], and audio [52]. Providing a comprehensive evaluation of these works may be an important future work. Second, this paper does not cover all possible scenarios of dataset usage in DL. We primarily evaluate the two most important and widely adopted settings: training from scratch (Section 4) and fine-tuning (Appendix H). Third, DATABench does not thoroughly explore finer-grained threat models, such as adversaries with varying levels of capabilities (an adversary with very limited computational resources or data). An exhaustive study of the attacker’s capabilities would be a meaningful direction for future work.
6 Conclusion
This paper presented a comprehensive adversarial evaluation of dataset auditing in Deep Learning (DL) systems, addressing the critical gap between existing dataset auditing methods and their resilience against dedicated attacks. We introduced a novel internal feature and external feature taxonomy to classify auditing techniques, formally defined evasion and forgery attack objectives, and proposed systematic attack strategies to realize them. These contributions led to the development of DATABench, a benchmark featuring 17 evasion and 5 forgery attacks. DATABench also provides an extensible toolbox to design new dataset auditing methods and attacks. This toolbox can also facilitate the unified evaluation of auditing methods. Our empirical experiments of 9 representative auditing methods against DATABench revealed some critical findings, $e.g.$, none of them demonstrated sufficient robustness or distinctiveness under adversarial settings. These results highlight the significant vulnerability of current dataset auditing methods and underscore the urgent need for developing novel and adversarially resilient dataset auditing methods.
Ethical Considerations
This paper focuses on dataset auditing in Deep Learning (DL). The objective of this paper is to advance the understanding of dataset auditing and evaluate the robustness and distinctiveness of existing methods from an adversarial perspective, which is critical to ensuring the transparency and trustworthiness of dataset auditing methods. Our paper intends to warn the community that existing dataset auditing methods may not be reliable under adversarial settings. All datasets used in our papers are publicly available. We also do not use any personally identifiable information or sensitive data.
Appendix A Detailed Introduction to Dataset Auditing implemented in DATABench
Shokri et al. (MIA) [77]. Shokri et al. proposed the first membership inference attack (MIA), leveraging models’ differential behavior ($e.g.$, higher confidence) on members versus non-members. The method trains multiple ‘shadow models’ to generate labeled prediction vectors (member/non-member), which then train a final classifier to distinguish target model members. We used 10 shadow models in our experiments.
Rapid [36]. Rapid is a SOTA MIA addressing inherent errors in difficulty calibration methods of reference-based attacks. It combines the original membership signal ($e.g.$, loss) with the calibrated score from reference models. Rapid trains a scoring model mapping both scores to a final membership prediction, learning to correct limitations of relying solely on calibration. We used 4 reference models in our experiments.
Dataset Inference (DI) [61]. DI assesses if a suspicious model shows higher prediction margins for the owner’s data points. It generates feature embeddings capturing this margin from the suspicious model using white-box or black-box methods. A confidence regressor, pre-trained on the owner’s model and data, scores these embeddings. Finally, a statistical hypothesis test aggregates scores to determine if the suspicious model used the owner’s dataset.
Data-use Auditing (DUA) [42]. DUA is a general dataset auditing framework. The owner creates two marked versions of each datum (preserving utility, maximizing difference), publishing one and hiding the other. Then, it uses a black-box Membership Inference (MI) score function and compares scores on the published versus hidden versions (contrastive MI). A sequential hypothesis test then determines if the published data was used for training.
DVBW [54]. DVBW proposes the first and a general framework for backdoor-based dataset auditing. Specifically, DVBW utilizes poison-label backdoor attacks for dataset auditing and presents a hypothesis-test-based verification method. In our experiments, the target label is set to 1.
UBW-P/C [51]. UBW is proposed for harmless and stealthy dataset copyright protection using untargeted backdoors. UBW has two variants, UBW-P and UBW-C. UBW-P utilizes an untargeted poison-label backdoor attack that changes the labels of the poisoned samples randomly. In contrast, UBW-C adopts bi-level optimization to craft clean-label ($i.e.$, the labels of the watermarked samples do not change) samples. Finally, UBW also utilizes a hypothesis test to complete dataset auditing.
ZeroMark [31]. ZeroMark is proposed for secure dataset auditing without disclosing watermarks. Inspired by the intrinsic property that boundary gradients of watermarked DNNs align with the watermark pattern, ZeroMark generates boundary versions of benign samples and calculates their boundary gradients under a label-only black-box setting. It then performs a hypothesis test based on the cosine similarity between the gradients and the secret watermark pattern to achieve auditing.
Domain Watermark (DW) [33]. DW is proposed for harmless dataset auditing by identifying a hardly-generalized domain. Instead of causing misclassification, DW ensures watermarked models correctly classify specific ‘hard’ samples misclassified by benign models. DW employs bi-level optimization to generate the domain and craft visually indistinguishable clean-label modified samples for stealth.
Appendix B Details of Evasion Attacks
Data Synthesis. Data synthesis involves generating artificial data, typically using generative models trained on the original dataset, to train the suspect model, thereby obscuring the direct link to the protected data. In our experiments, we use the DDIM diffusion model [37] as the generative model.
Denoising Filters. Denoising filters apply traditional image processing filters ($e.g.$, Gaussian, Median, Wavelet) to potentially remove noise-like external features (EF) embedded for auditing, effectively cleaning the data before model training. In our experiments, we set the kernel size to be 3.
Adversarial Training. Adversarial Training augments the training data with adversarially perturbed examples. We use two different implementations of adversarial training. AdvTraining-FGSM uses solely the FGSM method [29] to generate adversarial examples while AdvTraining-Hybrid adopts three methods, FGSM [29], PGD [59], and C&W [8].
Adaptive Splitting Poisoned Dataset (ASD) [28]. ASD is a training-time backdoor defense. ASD employs loss-guided splitting and meta-learning-inspired splitting to dynamically update two separate data pools. By leveraging the clean and poisoned data pools obtained through this adaptive partitioning, ASD effectively mitigates backdoors during training.
Noisy Output&Feature. Noise is injected into the model’s intermediate feature activations or final output predictions during inference to mask the subtle signals. In our experiments, we use Gaussian noise and the standard deviation is set to 0.5.
Reprogramming [13]. Reprogramming modifies the model’s inference process, often using learnable input and output transformations, to neutralize the specific abnormal behaviors. In our experiments, we use a 10-layer U-Net as the input transformation module and train the U-Net for 10 epochs.
SCALE-UP [32]. SCALE-UP is an effective black-box, input-level backdoor detection method. SCALE-UP identifies and filters malicious test samples by analyzing their prediction consistency during the pixel-wise amplification process.
Outlier Detection (OD). OD identifies input samples that are statistically different from typical data (potentially EF triggers), allowing the adversary to return a non-standard or random output to evade detection. In DATABench, we use two different models, $i.e.$, KNN and SVM, to achieve OD.
Appendix C Details of Forgery Attacks
Appendix D Other Detailed Experimental Settings
As default settings, we train ResNet-18 on CIFAR-10 from scratch for 90 epochs with a learning rate of 0.03. The batch size is set to 128. Moreover, since batch normalization in ResNet-18 does not support DP-SGD, we alter these layers with group normalization. For EF-based methods, we use the Blended method [12], which applies a weighed blend of a specific trigger pattern and the original image, with the trigger pattern assigned a weight of 0.2. The experiments are conducted using at most 4 NVIDIA RTX 3090 GPUs.
Appendix E Discussion of the Model Utility Degradation
To ensure a fair, simple, and reproducible evaluation, we adopt the classical or standard/default configurations for each attack implemented in our DATABench. While this choice may lead to some degradation in model utility under certain scenarios, we argue that the key findings of this paper remain valid for the following two main reasons.
  • Some evasion techniques are inherently designed with objectives that conflict with maximizing model utility. A representative example is DP-SGD, which deliberately injects noise into the training process to provide certified privacy guarantees. As demonstrated in our experiments, this mechanism substantially reduces test accuracy while effectively evading auditing methods. In such cases, the degradation in utility is an intentional and acceptable trade-off for achieving the desired level of privacy.
  • Arguably, a moderate reduction in a model’s final utility does not diminish the value gained from the unauthorized use of a dataset. Even if an adversary’s model attains lower accuracy after applying an evasion attack, the protected dataset has still been exploited and has contributed to the model’s training and development. Thus, the core concern of unauthorized data usage—whether in terms of privacy breaches or copyright infringements—remains unresolved.
Table 6: Results of the hybrid evasion attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Method Metric No Attack HA1 HA2 HA3
IF-based MIA Test Acc. (%, $\uparrow$) 87.58 78.75 56.93 74.34
Score ($\uparrow$) 0.62 0.00 0.00 0.00
Rapid Test Acc. ($\%,\uparrow$) 87.58 78.04 57.65 74.05
Score ($\uparrow$) 0.66 0.22 0.16 0.21
DI Test Acc. (%, $\uparrow$) 87.58 79.71 56.74 73.73
p-value ($\downarrow$) $10^{-25}$ 1.00 0.08 0.13
Diff ($\uparrow$) 1.33 0.00 0.00 0.00
DUA Test Acc. (%, $\uparrow$) 87.17 80.39 61.20 73.34
Cost (logits) (%, $\downarrow$) 16.66 99.77 100.00 22.17
Cost (label) (%, $\downarrow$) 85.97 99.86 100.00 23.77
EF-based DVBW Test Acc. (%, $\uparrow$) 86.58 78.10 55.69 70.40
p-value ($\downarrow$) 0.00 1.00 1.00 1.00
WSR (%, $\uparrow$) 100.00 11.47 100.00 19.36
UBW-P Test Acc. (%, $\uparrow$) 87.28 78.51 50.51 71.75
p-value ($\downarrow$) 0.00 1.00 1.00 1.00
WSR (%, $\uparrow$) 87.39 24.82 87.50 28.91
UBW-C Test Acc. (%, $\uparrow$) 83.05 72.81 57.17 70.55
p-value ($\downarrow$) $^{-122} 1.00 1$.00 1.00
WSR (%, $\uparrow$) 74.60 35.95 75.49 29.10
ZeroMark Test Acc. (%, $\uparrow$) 87.14 78.78 56.32 75.59
p-value ($\downarrow$) $^{-137} 1.00 1$.00 1.00
WSR (%, $\uparrow$) 100.00 13.88 100.00 25.03
DW Test Acc. (%, $\uparrow$) 86.87 77.11 63.61 42.95
WSR (%, $\uparrow$) 67.20 37.20 0.10 6.59
@@ Appendix E Discussion of the Model Utility Degradation
Appendix F Experiments of Hybrid Evasion Attacks
We also explore combining evasion attack methods from different phases to enhance overall attack effectiveness. Specifically, we hereby evaluate three types of hybrid evasion attacks: (1) HA1: AutoEncoder, Adversarial Training, and Reprogramming. (2) HA2: WaveletFilter, DP-SGD, and Randomized Smoothing. (3) HA3: WaveletFilter and Reprogramming. The results are shown in Table 6.
Takeaway 6: Combining attacks from different phases can significantly enhance the attack effectiveness.
Our empirical results corroborate this takeaway, demonstrating the heightened threat posed by hybrid evasion attacks. For instance, HA1 and HA3 successfully circumvent nearly all evaluated auditing methods, substantially degrading the performance metrics and causing audit failures. HA2 also demonstrates considerable efficacy, evading all the auditing methods, but also degrades the utility of the trained models. This finding highlights the feasibility of a sophisticated adversary to combine attacks to achieve better attack effectiveness.
Appendix G Experiments on Extra Models and Datasets
Experimental Settings. We conduct evaluations on the (100-class subset of) ImageNet dataset [17] using a different adversary’s model architecture, MobileViT [62], trained from scratch. We maintain the realistic assumption that if an auditor needed to train auxiliary models for their auditing method ($e.g.$, shadow models for MIA), they would use a standard architecture (ResNet-18), thus creating an architectural mismatch with the ViT suspect model.
Results. Consistent with our primary findings, the results in Table 7 show that existing dataset auditing methods remain highly vulnerable to evasion attacks in this more challenging setting. Both IF-based and EF-based auditing techniques frequently failed to correctly identify the usage of the ImageNet victim dataset when attacks like AutoEncoder-based denoising, adversarial training, or Reprogramming were employed. The architectural mismatch (transformer-based model and auditor’s CNN) and the increased dataset complexity do not confer robustness to the evaluated auditing methods.
Appendix H Evaluating Dataset Auditing in the Fine-tuning Scenario
@@
Experimental Settings. Beyond training models from scratch, fine-tuning pre-trained models on a task-specific dataset is a common and efficient paradigm in modern DL. We simulate the fine-tuning scenario using a ResNet-18 pre-trained on ImageNet-1000 and then fine-tune it using CIFAR-10 as the target (victim) dataset. Other experimental settings remain consistent with our main experiments in Section 4.

### v3

@@ 5.2 Experimental Settings
5.3 Results of Evasion Attacks
Tables 2-4 show the results of evasion attacks against ResNet-18 in the pre-processing, training, and post-training phases, respectively. Table 5 shows the results of representative evasion attacks against MobileViT [mehta2021mobilevit] trained on (a 100-class subset of) ImageNet [deng2009imagenet]. We maintain the realistic assumption that if an auditor needed to train auxiliary models for their auditing method ($e.g.$, shadow models for MIA), they would use a standard architecture (ResNet-18), thus creating an architectural mismatch with the ViT suspect model. We summarize our key findings as follows.
@@ 5.3 Results of Evasion Attacks
A key finding is that dataset auditing methods relying on the embedding of strong external features, exemplified by DVBW and UBW-P in our experiments, exhibit relatively higher robustness. Their robustness is particularly exhibited against challenging training-phase attacks like adversarial training and DP-SGD. The two methods are poison-label methods that necessitate changing the labels of part of the samples in the dataset. However, even these methods are not immune to evasion attacks, showing vulnerabilities to specific pre-processing data transformations ($e.g.$, WaveletFilter, AutoEncoder) and post-training inference manipulations ($e.g.$, Reprogramming). Besides, other EF-based methods tested using mild external features ($i.e.$, UBW-C, ZeroMark, DW) displayed considerably less robustness across various attack stages.
Moreover, IF-based methods generally exhibit moderate robustness, with DUA demonstrating relatively higher robustness. DUA enhances auditing effectiveness by injecting additional noise into the dataset. This mechanism improves robustness but also makes the audit results more susceptible to forgery attacks. We will further discuss the details in Section 5.5.
Takeaway 4: For modern models and large-scale datasets, the vulnerability of dataset auditing still holds.
Consistent with our primary findings, the results in Table 5 demonstrate that existing dataset auditing methods remain highly vulnerable to evasion attacks when auditing modern models (Vision Transformers) and large-scale datasets (ImageNet). Both IF-based and EF-based auditing techniques frequently failed to correctly identify the usage of the ImageNet victim dataset when attacks like AutoEncoder-based denoising, adversarial training, or Reprogramming were employed. The architectural mismatch (transformer-based model and auditor’s CNN) and the increased dataset complexity do not confer robustness to the evaluated auditing methods.
Table 7: Results of forgery attacks in the white-box scenario. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack FGSM PGD UAP TIFGSM VNIFGSM
IF-based MIA Score ($\downarrow$) 0.54 0.66 0.72 0.54 0.71 0.71
Rapid Score ($\downarrow$) 0.22 0.79 0.89 0.30 0.86 0.88
DI p-value ($\uparrow$) 0.01 $10^{-13}$ $10^{-5}$ $10^{-23}$ 0.97 0.01
Diff ($\downarrow$) 0.79 1.35 0.98 1.61 -0.81 0.09
DUA Cost (logits) (%, $\uparrow$) 100.00 0.13 0.13 0.18 0.13 0.24
Cost (label) (%, $\uparrow$) 100.00 0.13 0.13 0.24 0.13 0.24
EF-based DVBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 10.01 29.92 98.37 65.60 69.77 96.77
UBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 11.73 86.55 100.00 71.08 95.69 100.00
ZeroMark p-value ($\uparrow$) 1.00 1.00 1.00 0.99 0.01 1.00
WSR (%, $\downarrow$) 10.01 29.93 98.44 68.28 71.68 96.88
DW WSR (%, $\downarrow$) 11.73 67.00 89.40 66.37 73.72 89.36
@@
5.4 Evaluating Dataset Auditing in the Fine-tuning Scenario
@@
Beyond training models from scratch, fine-tuning pre-trained models on a task-specific dataset is a common and efficient paradigm in modern DL. We simulate the fine-tuning scenario using a ResNet-18 pre-trained on ImageNet-1000 and then fine-tune it using CIFAR-10 as the target (victim) dataset. Other experimental settings remain consistent with our main experiments in Section 5.3.